Swimming pool length and width are 20m and 10m what is the area of the pool in (cm )
1
Expert's answer
2019-10-14T09:33:11-0400

"S=20m\cdot10m =200m^2=200\cdot(100sm)^2=
200\cdot10000sm^2=2\cdot10^6 sm^2"

Answer: "S=2\cdot10^6 sm^2"